# Machine Learning for Synthetic Data Generation: A Review

### v6

@@ I Introduction
Paper
application
generative AI
DNN
dataset
MedGAN [16]
healthcare
GAN
MLP
MIMIC/Sutter (Electronic health record)
MMCGAN [17]
healthcare & CV
GAN
CNN
chest CT images
DeepSynth [18]
healthcare & CV
GAN
CNN
rat kidney tissue (microscope image)
ChemSpaceE [19]
drug
VAE
GNN
ZINC (drug molecule) [20]
JTVAE [21]
drug
VAE
GNN
ZINC (drug molecules) [20]
REINVENT [22]
drug
RL
RNN
ZINC (drug molecules) [20]
CORE [23]
drug
VAE
GNN
ZINC (drug molecule) [20]
RGA [24]
drug
RL
geometric NN
ZINC and TDC [25]
CorGAN [26]
healthcare
GAN
CNN
MIMIC-III dataset, UCI Epileptic Seizure Recognition dataset
DAAE [27]
healthcare
VAE+GAN
recurrent autoencoder
MIMIC-III, UT Physicians clinical databases
HAPNEST [28]
healthcare
approximate Bayesian computation (not deep learning)
NA (w.o. DNN)
Genomes Project and HGDP datasets
synthpop [29]
healthcare
proper synthesis
Statistical hypothesis testing
SD2011
CycleGAN [30]
vision
GAN
CNN
pix2pix
DP-CGAN [31]
vision
GAN
deep CGAN
MNIST
BigGANs [32]
vision
GAN
large scale GAN
ImageNet
VideoDiff [33]
vision
diffusion
CNN
BAIR Robot Pushing, Kinetics-600
VQ-VAE [34]
vision
VAE
PixelCNN
ImageNet
GIRAFFE [35]
vision
GAN
CNN
CompCars, LSUN Churches, and FFHQ
Wavegrad [36]
TTS
diffusion
gradient-based sampling
LJ Speech
TTS-GAN [37]
TTS
GAN
auto-regressive model
Tacotron2
Seq-GAN [38]
NLP
GAN+RL
CNN
Nottingham dataset
BLEURT [39]
NLP
Language model
BERT
WebNLG Competition dataset
TextGen-RL [40]
NLP
RL
LSTM
SynBench [41]
NLP
conditional Gaussian mixture
CIFAR10
RelGAN [42]
image and text
GAN
CNN
COCO Image Captions dataset
DPGM [43]
audio and text
generative artificial neural networks
differentially private kernel k-means
MNIST, anonymized Call Detail Record (CDR)
WaveGAN [44]
audio
GAN
DCGAN
Speech Commands Dataset
Wavenet [45]
audio
GAN
LSTM
CSTR voice corpula (multi-channel English audio)
Stutter-TTS [46]
audio
phonetic encoder and the decoder
CNN
recordings
Quant GANs [47]
business
GAN
MLP+ Temporal convolutional networks(TCN)
simulated data
CGAN [48]
business
GAN
CNN
Vector autoregressive (VAR) time series
PATE-GAN [49]
business
GAN
PRIVATE AGGREGATION OF TEACHER ENSEMBLES (PATE)
Kaggle
CollGAN [50]
physics (particle collision)
VAE/GAN
MLP
ATLAS
@@ IV Generative AI
Method
supervision
NN architecture
MLE
with latent variable
paper
Language model (LM)
no
autoregressive model
yes
no
[136]
self-supervised learning (SSL)
no
encoder (representation)
yes
no
variational autoencoder (VAE)
no
encoder-decoder
yes
yes
[137]
generative adversarial network (GAN)
no
generator & discriminator
yes
yes
[138, 139]
diffusion (score-based model) model
no
representation
yes
no
[140]
reinforcement learning (RL)
yes
policy network or Q-network
no
no
[141]
@@ V-A Privacy Risks in Data Synthesis
Paper
privacy-enhancing techniques
generative AI
DNN
data format
[165]
plausible deniability
probabilistical transform
NA (w.o. DNN)
attribute
[43]
differential privacy
generative artificial neural networks
kernel k-means
image and text
[166]
data replacement and item regularizer
latent space projection
MLP
attributes
[167]
differential privacy
GAN
CNN
image
[49]
differential privacy
GAN
Private Aggregation of Teacher Ensembles
attribute
[168]
differential privacy
GAN
CNN
EHR
[169]
differential privacy
GAN
CNN
image and EHR
[14]
differential privacy
autoencoder
autoencoder
attribute
[170]
differential privacy
MRF
NA (w.o. DNN)
attribute
[171]
compressive privacy
GAN
CNN
image
[172]
differential privacy
GAN
CNN
image
[173]
local differential privacy
Maximum Entropy estimation
NA (w.o. DNN)
attribute
[174]
differential privacy
Maximum Entropy estimation
NA (w.o. DNN)
attribute
[175]
differential privacy
GAN
CNN
image
[176]
differential privacy
Bayesian network
NA (w.o. DNN)
attribute
[177]
differential privacy
statistical database
NA (w.o. DNN)
attribute
[178]
differential privacy
Graduate Update Method
NA (w.o. DNN)
attribute
[179]
differential privacy
autoencoder
autoencoder
image and attribute
@@ VIII Evaluation Strategy
    Training on synthetic dataset and testing on the real dataset (TSTR). This strategy involves using synthetic data to train machine learning models and assessing their prediction performance on real test data in downstream applications. High performance on real test data indicates that the synthetic data has successfully captured essential characteristics of the real data, making it a useful proxy for training. For example, [207] employ synthetic data to train machine learning models and assess their prediction performance on real test data in downstream applications. TSTR can provide insights into the effectiveness of synthetic data for training machine learning models in a wide range of tasks and domains.
@@ X Conclusion
In conclusion, machine learning has revolutionized various industries by enabling intelligent computer systems to autonomously tackle tasks, manage and analyze massive volumes of data. However, machine learning faces several challenges, including data quality, data scarcity, and data governance. These challenges can be addressed through synthetic data generation, which involves the artificial annotation of information generated by computer algorithms or simulations. Synthetic data has been extensively utilized in various sectors due to its ability to bridge gaps, especially when real data is either unavailable or must be kept private due to privacy or compliance risks.

### v8

@@ I Introduction
Paper application generative AI DNN dataset
MedGAN [16] healthcare GAN MLP MIMIC/Sutter (Electronic health record)
MMCGAN [17] healthcare & CV GAN CNN chest CT images
DeepSynth [18] healthcare & CV GAN CNN rat kidney tissue (microscope image)
ChemSpaceE [19] drug VAE GNN ZINC (drug molecule) [20]
JTVAE [21] drug VAE GNN ZINC (drug molecules) [20]
REINVENT [22] drug RL RNN ZINC (drug molecules) [20]
CORE [23] drug VAE GNN ZINC (drug molecule) [20]
RGA [24] drug RL geometric NN ZINC and TDC [25]
CorGAN [26] healthcare GAN CNN MIMIC-III dataset, UCI Epileptic Seizure Recognition dataset
DAAE [27] healthcare VAE+GAN recurrent autoencoder MIMIC-III, UT Physicians clinical databases
HAPNEST [28] healthcare approximate Bayesian computation (not deep learning) NA (w.o. DNN) Genomes Project and HGDP datasets
synthpop [29] healthcare proper synthesis Statistical hypothesis testing SD2011
CycleGAN [30] vision GAN CNN pix2pix
DP-CGAN [31] vision GAN deep CGAN MNIST
BigGANs [32] vision GAN large scale GAN ImageNet
VideoDiff [33] vision diffusion CNN BAIR Robot Pushing, Kinetics-600
VQ-VAE [34] vision VAE PixelCNN ImageNet
GIRAFFE [35] vision GAN CNN CompCars, LSUN Churches, and FFHQ
Wavegrad [36] TTS diffusion gradient-based sampling LJ Speech
TTS-GAN [37] TTS GAN auto-regressive model Tacotron2
Seq-GAN [38] NLP GAN+RL CNN Nottingham dataset
BLEURT [39] NLP Language model BERT WebNLG Competition dataset
TextGen-RL [40] NLP RL LSTM
SynBench [41] NLP conditional Gaussian mixture CIFAR10
RelGAN [42] image and text GAN CNN COCO Image Captions dataset
DPGM [43] audio and text generative artificial neural networks differentially private kernel k-means MNIST, anonymized Call Detail Record (CDR)
WaveGAN [44] audio GAN DCGAN Speech Commands Dataset
Wavenet [45] audio GAN LSTM CSTR voice corpula (multi-channel English audio)
Stutter-TTS [46] audio phonetic encoder and the decoder CNN recordings
Quant GANs [47] business GAN MLP+ Temporal convolutional networks(TCN) simulated data
CGAN [48] business GAN CNN Vector autoregressive (VAR) time series
PATE-GAN [49] business GAN PRIVATE AGGREGATION OF TEACHER ENSEMBLES (PATE) Kaggle
CollGAN [50] physics (particle collision) VAE/GAN MLP ATLAS
@@ IV Generative AI
Method supervision NN architecture MLE with latent variable paper
Language model (LM) no autoregressive model yes no [136]
self-supervised learning (SSL) no encoder (representation) yes no
variational autoencoder (VAE) no encoder-decoder yes yes [137]
generative adversarial network (GAN) no generator & discriminator yes yes [138, 139]
diffusion (score-based model) model no representation yes no [140]
reinforcement learning (RL) yes policy network or Q-network no no [141]
@@ V-A Privacy Risks in Data Synthesis
Paper privacy-enhancing techniques generative AI DNN data format
[165] plausible deniability probabilistical transform NA (w.o. DNN) attribute
[43] differential privacy generative artificial neural networks kernel k-means image and text
[166] data replacement and item regularizer latent space projection MLP attributes
[167] differential privacy GAN CNN image
[49] differential privacy GAN Private Aggregation of Teacher Ensembles attribute
[168] differential privacy GAN CNN EHR
[169] differential privacy GAN CNN image and EHR
[14] differential privacy autoencoder autoencoder attribute
[170] differential privacy MRF NA (w.o. DNN) attribute
[171] compressive privacy GAN CNN image
[172] differential privacy GAN CNN image
[173] local differential privacy Maximum Entropy estimation NA (w.o. DNN) attribute
[174] differential privacy Maximum Entropy estimation NA (w.o. DNN) attribute
[175] differential privacy GAN CNN image
[176] differential privacy Bayesian network NA (w.o. DNN) attribute
[177] differential privacy statistical database NA (w.o. DNN) attribute
[178] differential privacy Graduate Update Method NA (w.o. DNN) attribute
[179] differential privacy autoencoder autoencoder image and attribute
@@ VIII Evaluation Strategy
    Training on synthetic dataset and testing on the real dataset (TSTR). This strategy involves using synthetic data to train machine learning models and assessing their prediction performance on real test data in downstream applications. High performance on real test data indicates that the synthetic data has successfully captured essential characteristics of the real data, making it a useful proxy for training. For example, [207] employs synthetic data to train machine learning models and assess their prediction performance on real test data in downstream applications. TSTR can provide insights into the effectiveness of synthetic data for training machine learning models in a wide range of tasks and domains.
@@ X Conclusion
In conclusion, machine learning has revolutionized various industries by enabling intelligent computer systems to autonomously tackle tasks, manage and analyze massive volumes of data. However, it still faces several challenges, including data quality, data scarcity, and data governance. These challenges can be addressed through synthetic data generation, which involves the artificial annotation of information generated by computer algorithms or simulations. Synthetic data has been extensively utilized in various sectors due to its ability to bridge gaps, especially when real data is either unavailable or must be kept private due to privacy or compliance risks.